OpenStudy (anonymous):

If we know f(3)=2.1, f'(3)=-3.6,f"(3)=5.4 and |f'''(z)|<9.2 for z between 3 and 3.012. What is the best esitmate of f(3.012) and what is the error bound"?" ;

OpenStudy (anonymous):

Sounds like taylor series\[f(x)=\frac{f(a)}{0!}(x-a)^0+\frac{f'(a)}{1!}(x-a)^1+\frac{f''(a)}{2!}(x-a)^2...\] With a=3

OpenStudy (anonymous):

I'm not sure how |f'''(z)|<9.2 would help with that, though.

OpenStudy (anonymous):

Then just insert x=3.012 for the approximation.

OpenStudy (anonymous):

expansion is maybe \[2.1-3.6(x-3)+2.7(x-3)^2+\frac{f'''(3)}{6}(x-3)^3+...\]

OpenStudy (anonymous):

last term gives the bound for the error if i am not mistaken.

OpenStudy (anonymous):

\[\frac{9.2}{6}(.012)^3\] would be the error i think

OpenStudy (anonymous):

are any of you familiar with maple? i need some help checking some work i did